OpenStudy (anonymous):

Find the value of sin θ from a right triangle with base of 9ft, vertical side 12 ft, and theta is the angle next to the right angle. /| / | 12 ft -- 9 ft the base angle to the left of the nine is theta.

OpenStudy (anonymous):

\[\tan \theta = \frac{12}{9} = \frac{4}{3} \] \[ h = \sqrt{4^2 + 3^2}\times 3 = 15\times3\] \[\sin\theta = \frac{12}{15}= \frac{4}{5}\]

OpenStudy (anonymous):

and the cotangent?

OpenStudy (anonymous):

\[\cot\theta = \frac{1}{\tan \theta} = \frac{3}{4}\] have a nice day :)

OpenStudy (anonymous):

sorry, again, i meant cosine lol

OpenStudy (anonymous):

\[\cos \theta = \frac{9}{15} = \frac{3}{5}\]